OpenStudy (anonymous):

Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school?

myininaya (myininaya):

b+g=94 (it said total of boys and girls is 94) b=22+2g (it said b=22 more than twice the number of girls)

myininaya (myininaya):

so take equation two in plug into equation one and solve for g then take that g put it in equation 2 to find b

myininaya (myininaya):

22+2g+g=94 <solve for g 22+3g=94 3g=72 g=24 now plug this into second equation to find b and presto-magic-o or whatever the saying is

myininaya (myininaya):

well you can plug into either equation whatever

OpenStudy (anonymous):

can you show the end cuz i dont get it